Two identical thermometers made of Pyrex glass contain, respectively, identical volumes of mercury and methyl alcohol. If the expansion of the glass is taken into account, calculate how many times greater is the distance between the degree marks on the methyl alcohol thermometer than the distance on the mercury thermometer


1
Expert's answer
2021-06-16T16:38:10-0400

"h-height;\\b-coefficient \ of \ volume \ expansion;\\
a- alcohol, \ g - glass, \ m - mercury; \\
\frac{h_a}{h_m}=\frac{b_a-b_g}{b_m-b_g}=\frac{1200 \times10^{-6 } - 9.9\times10^{-6 }}{182\times10^{-6 }-9.9\times10^{-6 }}= 6.915 \\
{Answer:} \ 6.915 \ times."